I need help with this word problem which needs to be solved using a system of two linear equations. Thank you in advance for your help.
A university bookstore recently sold a wirebound graph-paper notebook for $2.50, and a college-ruled notebook for $2.30. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $118.60. How many of each type were sold?

Answer by jojo14344(1513)

Sure!
First, Let's assigned "x" for wirebound notebk, and "y' for college notebk, and we know the total is 50. Equating the fact, "x+y=50" right? -------- eqn 1
Next, $2.50 of each wirebound were sold PLUS $2.30 each of college also sold and total to $118.60. Equating that fact,
$2.50(x) + $2.30(y) = $118.60, isn't it? ----------- eqn 2
Go back eqn 1 we get "y=50-x" and substitute this value of y in eqn 2. Continuing,
$2.50(x) + $2.30(50-x) = $118.60
$2.50(x) + $115 -$2.30(x) = $118.60
$0.20(x) = $118.60-$115
$0.20(x) = $ $3.60
x= 18, total pieces for wirebound notebook
in eqn 1 we get y=50-18= 32 pieces, total for college ruled notebk
In doubt? Go back eqn 2,
$2.50(18) + $2.30(32) =$118.60
$45 + $73.60 = $118.60
$118.60 = $118.60, cool!
Thank you,
Jojo